# Chemical enrichment of metal-poor stars orbiting massive black hole companions

### v2

@@ 3 Methods and Numerical Setup
To simulate the accretion onto a lighter companion, we set up an unequal mass binary in a wind tunnel. In this setup, the black hole-star binary is originally placed at rest at the origin of the grid, while a uniform wind is injected from the left boundary. Following A. Antoni et al. (2019), the binary’s orbit is initially circular and is allowed to evolve self-consistently as it accretes material (and the corresponding momentum) from the background gas and is subjected to drag forces. The center of mass of the system is not fixed at the origin, so the reaction of the binary to the gaseous medium may induce a center of mass motion.
@@
We use the FLASH hydrodynamics code (B. Fryxell et al., 2000) to solve the fluid equations over time. FLASH is a Eulerian code and we use the directionally split Piecewise Parabolic Method Riemann solver (P. Colella & P. R. Woodward, 1984). We construct a three-dimensional hydrodynamics setup where a uniform wind is injected with velocity $v_{\infty}$ from the x-x boundary in the +x+x direction. The +x+x, $\pm y$, and $\pm z$ boundaries are all outbound boundaries, while the x-x is an inbound boundary. The computational domain is $(-3,3)R_{a}\times(-3,3)R_{a}\times(-3,3)R_{a}$. The maximum and minimum refinement levels are set to 6 and 2, respectively, with a maximum cell size of $R_{a}/16$ and a minimum of $R_{a}/256$.
@@
The black hole-star binary is modeled by adding two active sinks (C. Federrath et al., 2010), which allows us to track the accretion and forces on both the black hole and the companion throughout the simulation. We initially place the center of mass of the system at the origin of the grid, $R_{\rm{CoM}}=0$ at time $t=0$. The center of mass of the binary is set to be at rest, $\bf{v}_{\rm{CoM}}=0$, and the particles are put in a circular orbit in the xyx-y plane with separation $a_{0}$ and angular momentum vector pointing in the +z+z direction. The sinks have a physical radius of $R_{s}=0.05R_{a}$, based on resolution studies, assuring that $R_{s}/\delta_{\rm{min}}=3$, where $\delta_{\rm{min}}$ is the length of the finest cell size (A. Antoni et al., 2019). The total mass of the system is fixed from one simulation to another, such that $M_{\rm{BH}}+M_{*}=M_{\rm{T}}$ where $M_{\rm{T}}$ is set to $(2G)^{-1}$ and $q={M_{*}}/{M_{\rm{BH}}}$. The simulations are carried out in dimensionless units where we set $v_{\infty}=1$ and $\rho_{\infty}=10$. This yields an accretion radius of $R_{a}=1$ which we use as the length scale of the problem. Time in the simulation has units of $R_{a}/v_{\infty}=1$.
@@
We investigate mass ratios of $q^{-1}=2,3,4,5$ for two distinct separations: $a_{0}=0.42$ and $1.0R_{a}$. Simulations with $a_{0}<0.42$ required significantly higher resolution and smaller sink radii to be reliably resolved (A. Antoni et al., 2019). For most of the simulations, we investigate cases where the gaseous medium has an adiabatic index of $\gamma=5/3$, except for the case where we keep $q$ and $a_{0}$ constant and vary $\gamma$. A $\gamma=1.1$ equation of state is used to simulate the properties of the flow when significant cooling occurs in the gas, which we denote as the quasi-isothermal case. The simulations with varying adiabatic index are presented in Appendix B. For all simulations in this work, the Mach number of the gas relative to the center of mass of the system is initially set to $\mathcal{M}=2.0$. The use of dimensionless parameters allows the results from these simulations to be scalable to a variety of mass ratio binaries across a wide range of astrophysical regimes, especially for mass ratios that are difficult to resolve numerically ($q^{-1}\gtrsim 5$).
@@ 4 Numerical Results
Figure 3: Density profiles from the center of mass of BH-star binaries for different mass ratios at a fixed initial separation of $a_{0}=0.42\,R_{a}$ (left) and $a_{0}=1.0\,R_{a}$ (right). From top to bottom, $q^{-1}=2,3,4,5$. The circles show the locations of the BHs and the star shaped symbols show the locations of the stars relative to the center of mass of the binary. We can see that the stars experience densities that are $\approx 6-10$ times larger than the background one. The gray dotted line shows the best fit to the single object between $0.1\text{ and }1.0R_{a}$, $\rho\propto r^{-0.81}$. Left inset: Orbit averaged accretion rate for the stellar companion compared to its mass accretion rate in isolation. This is plotted for different inverse mass ratios $q^{-1}$, at fixed orbital separation of $a_{0}=0.42$. The best fit line is plotted which yields the relation $\langle\dot{M}\rangle\propto\left(q^{-1}\right)^{1.15\pm 0.2}$.
@@ 4.2 Mass ratio and separation dependencies
We investigate how the local density experienced by the less massive companion in binary systems changes with separation for different mass ratios. Figure 3 shows the density distribution for two different initial separations $a_{0}/R_{a}$, each featuring four distinct mass ratios: $q^{-1}=2,3,4,5$. The uppermost black line labeled as $q^{-1}=0$ denotes the density profile for the case of a single object with total mass equal to that of the binary.
For both of the initial separations studied in Figure 3, the density redistribution approaches that of the single object as the mass ratio decreases ($q^{-1}$ increases). That is, for highly unequal mass ratios, the density redistribution profile will closely resemble the profile of a single mass. This is to be expected, as the smaller object has a diminishing influence and the majority of the redistribution of the gas is governed by the massive black hole. In these examples, where $a_{0}/R_{a}$ is held constant, we can see the trend where the physical separation between the objects increases as $q$ decreases. This trend is more pronounced in cases where the initial ratio $a_{0}/R_{a}$ is smaller, such as $0.42$ compared to $1.0$. In the single body scenario, the peak density reaches approximately 30 times the background density. In contrast, the secondary stars experience around 6 times the density enhancement for $a_{0}/R_{a}\approx 1.0$ and about 10 times the enhancement for $a_{0}/R_{a}\approx 0.42$.
@@
In the left inset of Figure 3, we plot the orbit averaged mass accretion rate as a function of $q^{-1}$. The mass accretion rate values are normalized to the mass accretion rate of the star if it was found in isolation in a medium with background density $\rho_{\infty}$. We derive $\langle\dot{M}\rangle\propto\left(q^{-1}\right)^{\xi}\propto\left(q^{-1}\right)^{1.15\pm 0.2}$. The best fit relation for $a_{0}=1.0$ is also consistent with $\xi\approx 1$. Here we directly probe the relation between the mass ratio and mass accretion rate presented in Eq. 9. We conclude that the density enhancement experienced by the companion is altered by the mass ratio as $\rho_{\rm{rel}}\propto(q^{-1})^{1.15\pm 0.2}\approx q^{-1}$. In high-density environments, cooling will further enhance the mass accretion rate (Appendix B). Thus, these relationships should be seen as offering a conservative limit.
@@ 5 Chemical enrichment of metal-poor stars orbiting Gaia BH3-like Black Holes
A recently discovered black hole orbiting a low-mass star was detected using astrometry with the Gaia satellite. This system consists of a black hole with a mass of $32.70\pm 0.82\,M_{\odot}$, which is orbited by a star with a mass of $0.76\pm 0.05\,M_{\odot}$ and a measured metallicity of ${\rm Fe/H}]=-2.56\pm 0.1$. The system has an orbital period of $4,235.1\pm 98.5$ days, a semi-major axis of $16.17\pm 0.27\,\text{AU}$, and an eccentricity of $0.7291\pm 0.0048$ (Gaia Collaboration et al., 2024). This system acts as a test bed for studying the accretion history of low metallicity stars that orbit massive black hole companions.
@@
We propose that the ISM pollutes low-mass companions through Bondi accretion. The ability of these companion stars to capture newly synthesized metals from the ISM is significantly enhanced by the gravity of their massive partners. In this scenario, the surface compositions of these companions would differ from their cores, with low-mass stars that have the thinnest convective layers experiencing the highest degree of pollution (i.e., less dilution). Low-mass stars, such as the companion of BH3, undergo mixing as they evolve away from the main sequence (A. I. Karakas & J. C. Lattanzio, 2014). The first mixing event is known as the first dredge-up, which occurs once the hydrogen in the core is depleted. As the core contracts, it leads to the expansion of the star and the deepening of its convective envelope. The first dredge-up submerges and dilutes any species that are only present on the stellar surface (R. G. Gratton et al., 2000; A. Olejak et al., 2020). This occurs, for instance, when a star accretes material from a disrupted planet (e.g., M. MacLeod et al., 2018; M. Soares-Furtado et al., 2021; R. Yarza et al., 2023). The resulting signatures in the star’s photosphere are influenced by internal mixing processes, which can weaken or alter the observed enrichment over time (M. Soares-Furtado et al., 2021).
@@
To quantify the observational signature of external pollution, we use, in what follows, Modules for Experimentation in Stellar Astrophysics (MESA; B. Paxton et al., 2011, 2013, 2015, 2018, 2019; A. S. Jermyn et al., 2023) and adopt the low_z test suite inlist. This setup facilitates the evolution of $0.8\,M_{\odot}$ low metallicity stars from the pre-main sequence through the asymptotic giant branch phase. The range of masses of the convective region during the main sequence is $1.47\times 10^{-3}\,M_{\odot}$ for stars with initial [Fe/H] =2=-2 and $1.07\times 10^{-3}\,M_{\odot}$ for [Fe/H] =5=-5. If the ISM contaminated the surface of the star during its main sequence phase, convection would have mixed this material throughout the convective envelope. When the star transitions off the main sequence, the first dredge-up would naturally decrease the surface abundance of the accreted elements (e.g., E. N. Kirby et al., 2020). Inspired by this, we examine the accretion history of low metallicity main-sequence stars orbiting massive black hole companions (Sections 5.1 and 5.2) and analyze how the resulting pollution signatures change as the companion stars evolve (Section 5.3). If this scenario is occurring in Gaia BH3, we would expect a depletion in the abundances of accreted metals at the surface of the companion star.
Figure 4: Top: Evolution of $a/R_{a}$ for a Gaia BH3-like binary. Namely, a $0.8\,M_{\odot}$ companion star orbiting around a $32.7\,M_{\odot}$ black hole at a separation of $16\,\rm{AU}$. The components used from Eris simulations are the velocity of the star particles relative to the ISM, which determines $R_{a}$, as well as the gas density as a function of $z$. The gray dotted line denotes where $a/R_{a}=1$, such that for all values under the line, the binary has a shared accretion radius. Bottom: The companion’s mass accretion history of iron from formation to today. In both panels, dark curves show the median values and the shaded region show the 6868% scatter about the median.
@@ 5.1 The mass accretion history of Gaia BH3-like binaries in the Milky Way
In order to calculate the amount of metals accreted over the companion star’s history, we use the stellar trajectories derived by S. Shen et al. (2017) from the Eris simulation (J. Guedes et al., 2011). These trajectories have information about the velocity of the system as well as the density and metallicity of the gas encountered by the stellar particles in the simulation. The Eris simulation is a high-resolution cosmological zoom-in simulation of a Milky Way analog from redshift $z=12$ to $z=0$. In the simulation, new star particles inherit the metallicity of their parent gas particles. These star particles return energy, mass, and metals to the ISM through core-collapse and type Ia supernovae (S. Shen et al., 2015). For further details about the Eris simulation, we refer the reader to J. Guedes et al. (2011), and for details on the stellar particle tracing we refer the reader to S. Shen et al. (2017).
@@
Following S. Shen et al. (2017), we calculated the total amount of Fe accreted onto the $0.8\,M_{\odot}$ companion star over its evolutionary history. This star is orbiting a $32.7\,M_{\odot}$ black hole at a separation of 16 AU mimicking the properties of Gaia BH3. Using the fits for $\xi$ and $\beta$ in Equation 10, we calculate the overall factor by which the relative density experienced by the star is augmented due to its black hole companion. That is, $\rho_{\rm{rel}}=\rho_{\infty}(M_{\rm{BH}}/M_{*})(R_{a}/a)\approx 41\rho_{\infty}(R_{a}/a)$. This allows us to calculate the accretion rate onto the low-mass orbiting companion for each of the particle trajectories outlined in S. Shen et al. (2017), with the relative velocity to the gas being calculated self-consistently. We utilize a uniform random sample of 2,000 metallicity-selected star particles, which were selected by S. Shen et al. (2017) from all the Eris star particles residing in the halo, to compute the metal accretion.
In the top panel of Figure 4, we illustrate the evolution of $(a/R_{a})$ of the binary system as a function of redshift. It is important to note that for values greater than $z\approx 4$, the median $a/R_{a}$ falls below the dotted line (i.e., $a/R_{a}\lesssim 1$). This indicates that the companion is within the accretion radius of the black hole, allowing for vast enhancements in the mass accretion rate. In the bottom panel of Figure 4, we show the corresponding mass accretion rate of iron material onto the low-mass companion. The majority of the accretion occurs at $z\gtrsim 4$, with a sudden drop after this. This drop is associated with the last major merger in the galaxy (S. Shen et al., 2017). To quantify the observational signature of external pollution, we integrate over the total amount of iron accreted using median value derived from the particle trajectories of halo stars (S. Shen et al., 2017) for redshifts where $a/R_{a}<1$. In most cases, the companion’s contribution to the total metal mass accreted is negligible when $a/R_{a}\gtrsim 1$, due to the drastic decrease in the accretion rate. Assuming an initial birth [Fe/H] value of 5-5, we find that the polluted surfaces of the stellar companion, while still on the main sequence, can reach values of approximately 2.2-2.2 (Section 5.3). This inspection indicates that the metallicity of non-evolved companion stars orbiting massive black holes may be significantly altered from their original birth metallicity. We will explore this idea further in Section 5.2. When a stellar companion evolves off the main sequence, as is the case with the Gaia BH3 companion, the initial dredge-up process can, in contrast, significantly decrease any excess surface abundance created from the accreted metals. In Section 5.3, we analyze the expected level of metal depletion in the surface abundance of the Gaia BH3 companion and discuss whether one might be able to derive meaningful constraints on its birth metallicity.
@@
Figure 5: The surface abundance resulting from accretion for a main-sequence star companion with $0.8\,M_{\odot}$, orbiting a $M_{\rm{BH}}$ black hole at a separation $a/R_{a}$. We assume the star has a birth [Fe/H]=5=-5. The surface abundance is progressively enhanced for stars residing at tighter separations with heavier black hole companions. The teal star-shaped marker denotes the conditions for the $Gaia$ BH3 system if it were in the main sequence and when it was actively accreting in the early universe (see Figure 4 where $a/R\approx 0.3$ for $z\lesssim 4$). Top: In black, we show the probability density function of LIGO-Virgo-Kagra black hole masses (R. Abbott et al., 2023b). In blue, we show the probability density function of the masses of the black hole component of low mass X-ray binaries (F. Fortin et al., 2024). Right: We show the probability density function of [Fe/H] of Gaia DR3 thick disc and halo stars (B. Kim & S. Lépine, 2021).

### v3

@@ 3 Methods and Numerical Setup
To simulate the accretion onto a lighter companion, we set up an unequal mass binary in a wind tunnel. In this setup, the black hole-star binary is originally placed at rest at the origin of the grid, while a uniform wind is injected from the left boundary. Following Antoni2019, the binary’s orbit is initially circular and is allowed to evolve self-consistently as it accretes material (and the corresponding momentum) from the background gas and is subjected to drag forces. The center of mass of the system is not fixed at the origin, so the reaction of the binary to the gaseous medium may induce a center of mass motion.
@@
We use the FLASH hydrodynamics code (Fryxell2000) to solve the fluid equations over time. FLASH is a Eulerian code and we use the directionally split Piecewise Parabolic Method Riemann solver (Colella_1984). We construct a three-dimensional hydrodynamics setup where a uniform wind is injected with velocity $v_{\infty}$ from the x-x boundary in the +x+x direction. The +x+x, $\pm y$, and $\pm z$ boundaries are all outbound boundaries, while the x-x is an inbound boundary. The computational domain is $(-3,3)R_{a}\times(-3,3)R_{a}\times(-3,3)R_{a}$. The maximum and minimum refinement levels are set to 6 and 2, respectively, with a maximum cell size of $R_{a}/16$ and a minimum of $R_{a}/256$.
@@
The black hole-star binary is modeled by adding two active sinks (Federrath_2010), which allows us to track the accretion and forces on both the black hole and the companion throughout the simulation. We initially place the center of mass of the system at the origin of the grid, $R_{\rm{CoM}}=0$ at time $t=0$. The center of mass of the binary is set to be at rest, $\bf{v}_{\rm{CoM}}=0$, and the particles are put in a circular orbit in the xyx-y plane with separation $a$ and angular momentum vector pointing in the +z+z direction. The sinks have a physical radius of $R_{s}=0.05R_{a}$, based on resolution studies, assuring that $R_{s}/\delta_{\rm{min}}=3$, where $\delta_{\rm{min}}$ is the length of the finest cell size (Antoni2019). The total mass of the system is fixed from one simulation to another, such that $M_{\rm{BH}}+M_{*}=M_{\rm{T}}$ where $M_{\rm{T}}$ is set to $(2G)^{-1}$ and $q={M_{*}}/{M_{\rm{BH}}}$. The simulations are carried out in dimensionless units where we set $v_{\infty}=1$ and $\rho_{\infty}=10$. This yields an accretion radius of $R_{a}=1$ which we use as the length scale of the problem. Time in the simulation has units of $R_{a}/v_{\infty}=1$.
@@
We investigate mass ratios of $q^{-1}=2,3,4,5$ for two distinct separations: $a=0.42$ and $1.0R_{a}$. Simulations with $a<0.42$ required significantly higher resolution and smaller sink radii to be reliably resolved (Antoni2019). For most of the simulations, we investigate cases where the gaseous medium has an adiabatic index of $\gamma=5/3$, except for the case where we keep $q$ and $a$ constant and vary $\gamma$. A $\gamma=1.1$ equation of state is used to simulate the properties of the flow when significant cooling occurs in the gas, which we denote as the quasi-isothermal case. The simulations with varying adiabatic index are presented in Appendix B. For all simulations in this work, the Mach number of the gas relative to the center of mass of the system is initially set to $\mathcal{M}=2.0$. The use of dimensionless parameters allows the results from these simulations to be scalable to a variety of mass ratio binaries across a wide range of astrophysical regimes, especially for mass ratios that are difficult to resolve numerically ($q^{-1}\gtrsim 5$).
@@ 4 Numerical Results
Figure 3: Density profiles from the center of mass of BH-star binaries for different mass ratios at a fixed initial separation of $a=0.42\,R_{a}$ (left) and $a=1.0\,R_{a}$ (right). From top to bottom, $q^{-1}=2,3,4,5$. The circles show the locations of the BHs and the star shaped symbols show the locations of the stars relative to the center of mass of the binary. We can see that the stars experience densities that are $\approx 6-10$ times larger than the background one. The gray dotted line shows the best fit to the single object between $0.1\text{ and }1.0R_{a}$, $\rho\propto r^{-0.81}$. Left inset: Orbit averaged accretion rate for the stellar companion compared to its mass accretion rate in isolation. This is plotted for different inverse mass ratios $q^{-1}$, at fixed orbital separation of $a=0.42$. The best fit line is plotted which yields the relation $\langle\dot{M}\rangle\propto\left(q^{-1}\right)^{1.15\pm 0.2}$.
@@ 4.2 Mass ratio and separation dependencies
We investigate how the local density experienced by the less massive companion in binary systems changes with separation for different mass ratios. Figure 3 shows the density distribution for two different initial separations $a/R_{a}$, each featuring four distinct mass ratios: $q^{-1}=2,3,4,5$. The uppermost black line labeled as $q^{-1}=0$ denotes the density profile for the case of a single object with total mass equal to that of the binary.
For both of the initial separations studied in Figure 3, the density redistribution approaches that of the single object as the mass ratio decreases ($q^{-1}$ increases). That is, for highly unequal mass ratios, the density redistribution profile will closely resemble the profile of a single mass. This is to be expected, as the smaller object has a diminishing influence and the majority of the redistribution of the gas is governed by the massive black hole. In these examples, where $a/R_{a}$ is held constant, we can see the trend where the physical separation between the objects increases as $q$ decreases. This trend is more pronounced in cases where the initial ratio $a/R_{a}$ is smaller, such as $0.42$ compared to $1.0$. In the single body scenario, the peak density reaches approximately 30 times the background density. In contrast, the secondary stars experience around 6 times the density enhancement for $a/R_{a}\approx 1.0$ and about 10 times the enhancement for $a/R_{a}\approx 0.42$.
@@
In the left inset of Figure 3, we plot the orbit averaged mass accretion rate as a function of $q^{-1}$. The mass accretion rate values are normalized to the mass accretion rate of the star if it was found in isolation in a medium with background density $\rho_{\infty}$. We derive $\langle\dot{M}\rangle\propto\left(q^{-1}\right)^{\xi}\propto\left(q^{-1}\right)^{1.15\pm 0.2}$. The best fit relation for $a=1.0$ is also consistent with $\xi\approx 1$. Here we directly probe the relation between the mass ratio and mass accretion rate presented in Eq. 9. We conclude that the density enhancement experienced by the companion is altered by the mass ratio as $\rho_{\rm{rel}}\propto(q^{-1})^{1.15\pm 0.2}\approx q^{-1}$. In high-density environments, cooling will further enhance the mass accretion rate (Appendix B). Thus, these relationships should be seen as offering a conservative limit.
@@ 5 Chemical enrichment of metal-poor stars orbiting Gaia BH3-like Black Holes
A recently discovered black hole orbiting a low-mass star was detected using astrometry with the Gaia satellite. This system consists of a black hole with a mass of $32.70\pm 0.82\,M_{\odot}$, which is orbited by a star with a mass of $0.76\pm 0.05\,M_{\odot}$ and a measured metallicity of ${\rm Fe/H}]=-2.56\pm 0.1$. The system has an orbital period of $4,235.1\pm 98.5$ days, a semi-major axis of $16.17\pm 0.27\,\text{AU}$, and an eccentricity of $0.7291\pm 0.0048$ (gaia_bh3). This system acts as a test bed for studying the accretion history of low metallicity stars that orbit massive black hole companions.
@@
We propose that the ISM pollutes low-mass companions through Bondi accretion. The ability of these companion stars to capture newly synthesized metals from the ISM is significantly enhanced by the gravity of their massive partners. In this scenario, the surface compositions of these companions would differ from their cores, with low-mass stars that have the thinnest convective layers experiencing the highest degree of pollution (i.e., less dilution). Low-mass stars, such as the companion of BH3, undergo mixing as they evolve away from the main sequence (2014PASA...31...30K). The first mixing event is known as the first dredge-up, which occurs once the hydrogen in the core is depleted. As the core contracts, it leads to the expansion of the star and the deepening of its convective envelope. The first dredge-up submerges and dilutes any species that are only present on the stellar surface (2000A&A...354..169G; 2020A&A...638A..94O). This occurs, for instance, when a star accretes material from a disrupted planet (e.g., 2018ApJ...853L...1M; 2021AJ....162..273S; 2023ApJ...954..176Y). The resulting signatures in the star’s photosphere are influenced by internal mixing processes, which can weaken or alter the observed enrichment over time (2021AJ....162..273S).
@@
To quantify the observational signature of external pollution, we use, in what follows, Modules for Experimentation in Stellar Astrophysics (MESA; Paxton2011; Paxton2013; Paxton2015; Paxton2018; Paxton2019; Jermyn2023) and adopt the low_z test suite inlist. This setup facilitates the evolution of $0.8\,M_{\odot}$ low metallicity stars from the pre-main sequence through the asymptotic giant branch phase. The range of masses of the convective region during the main sequence is $1.47\times 10^{-3}\,M_{\odot}$ for stars with initial [Fe/H] =2=-2 and $1.07\times 10^{-3}\,M_{\odot}$ for [Fe/H] =5=-5. If the ISM contaminated the surface of the star during its main sequence phase, convection would have mixed this material throughout the convective envelope. When the star transitions off the main sequence, the first dredge-up would naturally decrease the surface abundance of the accreted elements (e.g., 2020ApJ...891L..13K). Inspired by this, we examine the accretion history of low metallicity main-sequence stars orbiting massive black hole companions (Sections 5.1 and 5.2) and analyze how the resulting pollution signatures change as the companion stars evolve (Section 5.3). If this scenario is occurring in Gaia BH3, we would expect a depletion in the abundances of accreted metals at the surface of the companion star.
@@ 5.1 The mass accretion history of Gaia BH3-like binaries in the Milky Way
In order to calculate the amount of metals accreted over the companion star’s history, we use the stellar trajectories derived by Shen_2017 from the Eris simulation (Guedes_2011). These trajectories have information about the velocity of the system as well as the density and metallicity of the gas encountered by the stellar particles in the simulation. The Eris simulation is a high-resolution cosmological zoom-in simulation of a Milky Way analog from redshift $z=12$ to $z=0$. In the simulation, new star particles inherit the metallicity of their parent gas particles. These star particles return energy, mass, and metals to the ISM through core-collapse and type Ia supernovae (2015ApJ...807..115S). For further details about the Eris simulation, we refer the reader to Guedes_2011, and for details on the stellar particle tracing we refer the reader to Shen_2017.
Following Shen_2017, we calculated the total amount of Fe accreted onto the $0.8\,M_{\odot}$ companion star over its evolutionary history. This star is orbiting a $32.7\,M_{\odot}$ black hole at a separation of 16 AU mimicking the properties of Gaia BH3. Given its Galactic coordinates, Gaia BH3 is classified as stellar system belonging to the disk in the Eris simulation (Pillepich_2015), where approximately 77% of stars formed in situ (i.e., within the Milky Way). In-situ stars are expected to experience higher accretion rates over their history, as the Galactic star-forming environment is typically denser and more metal-rich than for stars residing in satellite galaxies.
@@
Using the fits for $\xi$ and $\beta$ in Equation 10, we calculate the overall factor by which the relative density experienced by the star is augmented due to its black hole companion. That is, $\rho_{\rm{rel}}=\rho_{\infty}(M_{\rm{BH}}/M_{*})(R_{a}/a)\approx 40\rho_{\infty}(R_{a}/a)$. This allows us to calculate the accretion rate onto the low-mass orbiting companion for each of the particle trajectories outlined in Shen_2017, with the relative velocity to the gas being calculated self-consistently. We utilize a uniform random sample of 2,000 metallicity-selected star particles, which were selected by Shen_2017 from all the Eris star particles residing in the halo, to compute the metal accretion.
Figure 4: Top: Evolution of $a/R_{a}$ for a Gaia BH3-like binary. Namely, a $0.8\,M_{\odot}$ companion star orbiting around a $32.7\,M_{\odot}$ black hole at a separation of $16\,\rm{AU}$. The components used from Eris simulations are the velocity of the star particles relative to the ISM, which determines $R_{a}$, as well as the gas density as a function of $z$. The gray dotted line denotes where $a/R_{a}=1$, such that for all values under the line, the binary has a shared accretion radius. Bottom: Distribution of the companion’s cumulative iron accretion history from formation to the present day, derived from the same set of trajectories and accounting for the spread in relative velocities. In both panels, dark curves show the median values and the shaded region show the 6868% scatter about the median.
We calculate the evolution of $a/R_{a}$ and the corresponding iron accretion onto the low-mass companion in several steps. First, we compute $a/R_{a}$ for each stellar particle using its relative velocity. In the top panel of Figure 4, we show the evolution of $a/R_{a}$ for the binary system as a function of redshift $z$. Each binary experiences varying background densities over its evolutionary history. We use the local density, $\rho_{\infty}$, together with $a/R_{a}$ to scale the accretion rate for each individual trajectory. Second, we integrate the accreted iron along each trajectory to obtain the mass accretion history of the companion. The bottom panel of Figure 4 shows the resulting iron accretion rate as a function of redshift. The majority of the accretion occurs at $z\gtrsim 4$, with a sudden drop after this. This drop is associated with the last major merger in the galaxy (Shen_2017).
To quantify the observational signature of external pollution, we integrate the total amount of iron accreted using values derived from the particle trajectories of halo stars (Shen_2017) and assume an initial birth [Fe/H] of 5-5 to calculate the observed metallicity of the polluted surface of the stellar companion. For redshifts where $a/R_{a}<1$, we apply a multiplicative factor to account for enhanced accretion, while for redshifts where $a/R_{a}\gtrsim 1$, we assume isolated accretion. In most cases, the companion’s contribution to the total metal mass accreted is negligible when $a/R_{a}\gtrsim 1$, due to the sharp decline in the accretion rate.
There is a range of outcomes that can be obtained by integrating over all the particle trajectories. To illustrate how the results vary, we perform two calculations. In the first, we use the mean value of $a/R_{a}$ across all trajectories and scale the accretion rate using the mean background density, $\rho_{\infty}$. In this case, the metallicity of the polluted surface is [Fe/H] =3.4=-3.4. We also explore the impact of variations in the background density by considering cases where $\rho_{\infty}$ is shifted by $\pm 1\sigma$, which captures the typical spread in the environment experienced by the particles. With these variations, the metallicity of the polluted surface ranges from [Fe/H] =4.5=-4.5 to 2.3-2.3. We then repeat the same exercise, but instead of varying the background density, we use the mean background density of all particles and examine the impact of variations in $R_{a}/a$ on the results, which depend on the evolution of $R_{a}$ as a function of the binary’s velocity. We find surface [Fe/H] values ranging from 4.0-4.0 to 2.11-2.11, with a median value of 3.4-3.4 for a stellar companion on the main sequence.
Given the current $Gaia$ BH3’s Galactic coordinates, we track the evolution of representative stellar particles with similar properties at the end of the Eris simulations (i.e., $z=0$). These systems allow us to estimate the surface metallicity of the stellar companion to be approximately [Fe/H] =2.2=-2.2, assuming it were still on the main sequence (Section 5.3).
This analysis indicates that the metallicity of non-evolved companion stars orbiting massive black holes may be significantly altered from their original birth values. We explore this idea further in Section 5.2. When a stellar companion evolves off the main sequence, as is the case for the Gaia BH3 companion, the initial dredge-up process can substantially reduce any excess surface abundance produced by accreted metals. In Section 5.3, we analyze the expected level of metal depletion in the surface abundance of the Gaia BH3 companion and discuss whether it is possible to place meaningful constraints on its birth metallicity. We note that our simulations do not include radiative feedback from the black hole onto the surrounding gas. While this effect is not expected to strongly alter the metallicity results presented here, we provide an estimate of its impact in Appendix C.
@@
Figure 5: The surface abundance resulting from accretion for a main-sequence star companion with $0.8\,M_{\odot}$, orbiting a $M_{\rm{BH}}$ black hole at a separation $a/R_{a}$. We assume the star has a birth [Fe/H]=5=-5. The top panel shows the corresponding surface metallicity [Fe/H] calculated using the accretion history for stellar particles with Galactic locations similar to the Gaia BH3 system. The teal star-shaped marker denotes the conditions for the $Gaia$ BH3 system if it were in the main sequence and when it was actively accreting in the early universe (see Figure 4 where $a/R\approx 0.3$ for $z\lesssim 4$). The bottom panel calculates the accretion history of the binary using the mean values of $a/R_{a}$ across all stellar particle trajectories and scaling the accretion rate with the mean background density, $\rho_{\infty}$ (Figure 4). For both cases, the surface abundance is progressively enhanced for stars residing at tighter separations with heavier black hole companions. Top panel distribution: In black, we show the probability density function of LIGO-Virgo-Kagra black hole masses (LIGO). In blue, we show the probability density function of the masses of the black hole component of low mass X-ray binaries (Fortin_2024). Right panel distribution: We show the probability density function of [Fe/H] of Gaia DR3 thick disc and halo stars (Kim_2021).